# Bongard-RWR+: Real-World Representations of Fine-Grained Concepts in Bongard Problems

### v1

@@ 4 Experiments
We evaluate $4$ state-of-the-art open-access VLMs on Bongard-RWR+: InternVL2.5 78B (IVL2.5) chen2024expanding ; chen2024internvl , Qwen2-VL 72B-Instruct (Q2VL) bai2023qwen ; wang2024qwen2 , LLaVA-Next 110B (LLaVA) jiang2023mistral ; liu2024improved ; liu2024visual , and MiniCPM-o 2.6 8B (MCPM) yao2024minicpm ; yu2024rlaif . Main experiments are conducted on the largest available model versions. Smaller variants are covered in ablation studies. All models are evaluated at a fixed decoding temperature of $0.5$, using structured JSON output format enforced via the Outlines decoding backend willard2023efficient . Inference was performed on an internal computing cluster equipped with NVIDIA DGX A100 and H100 nodes. The largest models required up to $24$ hours of inference per task using $4$ H100 GPUs.
We developed a Similarity Classifier (SC) to serve as a non-parametric baseline. For image-based setups (I1S and I2S), it computes ViT-L/14 embeddings dosovitskiy2021an for all matrix images and the test image. For text-based tasks (D1S and D2S), embeddings of image descriptions are obtained with fine-tuned MiniLM minilm2025hugging ; wang2020minilm . Next, for each matrix side, the embedding that is most distant from the test embedding (based on Euclidean distance) is identified. The test input is then assigned to the side with the lower maximum distance (i.e., higher similarity). Additional details are provided in Appendix B.
@@
To better characterize model behavior, we categorized Bongard-RWR+ matrix concepts into $9$ semantic groups reflecting the key factor behind each concept: Size, Position, Count, Branching, Similarity, Contour, Shape, Rotation, and Angle. This grouping follows recent efforts advocating for semantically grounded evaluations to isolate model strengths and weaknesses mitchell2021abstraction ; odouard2022evaluating . The results are shown in Fig. 6. For $K=2$, all models exhibit broadly similar profiles across concept groups. However, for more challenging setups ($K>2$), differences emerge. For instance, InternVL2.5 retains high performance on Shape, Size, and Branching for $K=16$, with accuracy near $75\%$. These categories involve high-level visual attributes that VLMs appear to capture reliably. In contrast, InternVL2.5 accuracy drops below $50\%$ for Contour, Rotation, and Angle. These concepts often rely on subtle visual cues (Contour) or precise spatial relationships (Rotation and Angle), which current models struggle to capture reliably. Overall, performance patterns remain consistent across values of $K$, supporting the utility of such analysis as a diagnostic tool for characterizing model capabilities.
@@
Given the difficulty of the introduced dataset, we hypothesized that decomposing the problem via an intermediate captioning step may improve model performance. To test this, we generated image descriptions using each of the $4$ VLMs and used these captions as input to the same set of models for solving the side prediction task. Since the prediction task is purely text-based, we also evaluated the DeepSeek-R1 70B (DS-R1) reasoning model guo2025deepseek , which does not support visual input ("N/A" for I1S and I2S in Table 2) but offers strong language understanding. As shown in Fig. 7, final accuracy in the D1S task varies depending on both the captioning (rows) and prediction (columns) models. Descriptions produced by InternVL2.5 yield the highest overall accuracy, with Qwen2-VL also producing relatively strong captions. Interestingly, captions from MiniCPM-o 2.6 lead to better results than those from LLaVA-Next, despite MiniCPM-o 2.6 being much smaller, highlighting its promising image-to-text capabilities. Among prediction models (columns), Qwen2-VL achieves the best average performance, followed by InternVL2.5, indicating relatively strong text-based reasoning abilities of both models. DeepSeek-R1, while known for strong performance on standard benchmarks, ranks third, revealing potential limitations in reasoning over image descriptions. Using captions generated by the best-performing model, InternVL2.5, we report D1S and D2S task performance in Table 2. Compared to I1S and I2S, most models perform above chance in these settings, confirming that the intermediate captioning step helps models better ground their predictions.
@@
Figure 8: Concept Selection ablations. (a) Accuracy of models of varying sizes for $K=16$. (b) The impact of using grayscale images for $K\in\{2,4,8,16\}$. Differences w.r.t. color images are shown using lighter colors and are annotated in parentheses above the corresponding bars.
@@
Prior works have shown that VLM performance tends to scale with model size hoffmann2022training ; kaplan2020scaling . To examine whether this observation holds for AVR, we evaluated smaller variants of several models in the CS task on Bongard-RWR+. Specifically, we additionally tested 8B, 26B, and 34B variants of InternVL2.5, 7B, 32B, and 72B versions of LLaVA-Next and 7B variant of Qwen2-VL. We excluded additional MiniCPM-o 2.6 variants, as the 8B model already showed weak performance and no larger variants are currently available. As shown in Fig. 8(a), accuracy generally increases with model size. Among the smallest configurations (7B/8B), InternVL2.5 performs best, followed by Qwen2-VL, mirroring the relation observed for their largest versions (70+B). Notably, both InternVL2.5 8B and Qwen2-VL 7B outperformed MiniCPM-o 2.6 8B, underscoring the generally stronger AVR capacity of these two model families.
@@
To assess the role of color in abstract reasoning, we evaluated all $4$ VLMs on Bongard-RWR+/GS, using the CS task. As shown in Fig. 8(b), model performance remains comparable, or even improves, on grayscale images. For instance, LLaVA-Next consistently achieves higher accuracy across all $K$ (e.g. +5+5 p.p. for $K=2$). This outcome aligns with the dataset’s design – the underlying concepts are derived from structural properties of black-and-white synthetic BPs, where color is not semantically relevant. In this context, color may act as a visual distractor that adds additional complexity, especially for lower-performing models.
Figure 9: Functional equivalence of generated images and real-world images. Performance trends in the CS task ($K\in\{2,4,8,16\}$) on Bongard-RWR and Bongard-RWR+/LP for $P=2,3$.
@@
Since Bongard-RWR+ relies on generated images, a key question is whether such images are as effective as real ones for evaluating visual reasoning. To investigate this, we compared model performance on the introduced dataset with that on Bongard-RWR, which consists exclusively of real-world images. However, given that Bongard-RWR includes only $60$ instances, we constructed matrices based on smaller subsets of images per side to expand the dataset. For each BP in Bongard-RWR, we considered all pairs of $P$-element subsets from both matrix sides (analogously as described in Section 3.1), yielding ${6\choose P}^{2}$ new matrices per BP, i.e., $36$, $225$, $400$, $225$, $36$ matrices for $P=1,\ldots,5$, resp. We applied this method to $54$ Bongard-RWR BPs that were used to generate Bongard-RWR+, and then selected $P=2,3$ to ensure sufficient dataset size, and capped the number of sampled matrices per source BP at $100$, resulting in 5 4005\,400 matrices per $P$. We evaluated all $4$ VLMs in the CS task on these Bongard-RWR variants and compared the results to those on Bongard-RWR+/LP, which uses analogous matrix construction approach but with generated images. As shown in Fig. 9, model performance follows similar trends on both datasets – accuracy decreases consistently as $K$ increases, regardless of whether the images are real or generated. This supports the validity of our generation-based approach for capturing the challenges of AVR.
@@
As $P$ increases, models are presented with more demonstrations of the underlying concept. While this could enhance concept identification, it also increases the amount of visual information to process. Interestingly, Fig. 9 shows that InternVL2.5 and Qwen2-VL benefit from larger $P$ in the CS task, e.g., InternVL2.5’s accuracy improves by $2,3,6,8$ p.p. for $K=2,\ldots,5$, resp. when comparing the results on Bongard-RWR+/L2 with Bongard-RWR+/L3. In contrast, the performance of LLaVA-Next and MiniCPM-o 2.6 shows no consistent improvement, suggesting these models may struggle to leverage additional examples.
To explore this further, we evaluated models on Bongard-RWR+/LP with $P=2,\ldots,6$ using the I1S and D1S tasks. In I1S (Table 3(a)), both InternVL2.5 and Qwen2-VL show a mild upward trend with increasing $P$, with Qwen2-VL achieving the best accuracy of $57\%$ at $P=6$, and InternVL2.5 close behind at $55\%$. In D1S (Table 3(b)), accuracy generally improves with $P$, showing that the models can effectively utilize additional demonstrations for concept identification in a text-based setting. Notably, DeepSeek-R1 outperforms others for every $P>2$ (and performs on-par with InternVL2.5 for $P=2$), demonstrating its strength in this setting. Additionally, D1S scores are consistently higher than I1S scores, showing that the employed models benefit from the explicit image captioning step. Nevertheless, performance remains modest across both tasks, in I1S being clearly inferior to the SC baseline, which highlights a persistent gap in AVR among the contemporary models.
Finally, model accuracy on Bongard-RWR+/LP exceeds that on Bongard-RWR+ (e.g., InternVL2.5 scores $50\%$ and $57\%$ in I1S and D1S, resp. on Bongard-RWR+ (cf. Table 2), but reaches $55\%$ and $67\%$ on Bongard-RWR+/L6). This gap stems from the differences in dataset construction, which we analyze in detail in Appendix B.
@@
We also assessed the ability of the models to generate free-form descriptions of underlying concepts in the CG task. Models were prompted to describe the concept behind each matrix, using either raw image inputs or InternVL2.5’s image captions. The generated responses were evaluated using standard NLP metrics: BLEU papineni2002bleu , METEOR banerjee2005meteor , $\text{ROUGE}_{L}$ lin2004rouge , CIDEr lin2004rouge , and BERTScore zhang2020bertscore . The results, reported in Appendix B, show that the models consistently achieved low scores on all metrics, with no clear best-performing model. Manual inspection of selected outputs further confirmed that current models generally struggle to articulate the BP concepts, highlighting this task as a particularly challenging direction for future research.
@@ 5 Conclusions
We introduced Bongard-RWR+, a large-scale dataset for evaluating AVR capabilities using generated real-world-like images that depict abstract concepts originating from classical BPs. The dataset supports a diverse set of BP task formulations, including multiclass concept selection, binary side classification, and free-form text generation. Our results show that while current VLMs demonstrate certain ability to identify coarse-grained concepts, they consistently struggle with fine-grained concept recognition. Although explicit image captioning and increased visual demonstrations can improve performance, even the strongest models fall short in more demanding setups, in particular, the unconstrained textual answer generation.
@@
Our dataset generation pipeline still requires human oversight to ensure adherence to intended concepts; advancing generative modeling is needed to improve scalability. Our evaluation includes selected VLMs; broader community involvement is essential to benchmark additional methods on Bongard-RWR+. A systematic review of AVR performance across datasets, such as VCog-Bench cao2024visual , MindSet biscione2024mindset , or WAIS-IV galatzer2024cognitive , could further contextualize model capabilities. We elaborate on these promising research directions in Appendix C.
Appendix A Dataset variants

### v2

@@ 4 Experiments
We evaluate $4$ state-of-the-art open-access VLMs on Bongard-RWR+: InternVL2.5 78B (IVL2.5) (Chen et al., 2024a; b), Qwen2-VL 72B-Instruct (Q2VL) (Bai et al., 2023; Wang et al., 2024), LLaVA-Next 110B (LLaVA) (Jiang et al., 2023; Liu et al., 2024a; b), and MiniCPM-o 2.6 8B (MCPM) (Yao et al., 2025; Yu et al., 2025). Main experiments use the largest available models, with smaller variants covered in ablations. All models are evaluated at a fixed decoding temperature of $0.5$, using structured JSON output format enforced via the Outlines decoding backend (Willard and Louf, 2023). Inference was performed on an internal computing cluster equipped with NVIDIA DGX A100 and H100 nodes. The largest models required up to $24$ hours of inference per task using $4$ H100 GPUs.
We developed a Similarity Classifier (SC) to serve as a non-parametric baseline. For image-based setups (I1S and I2S), it computes ViT-L/14 embeddings (Dosovitskiy et al., 2021) for all matrix images and the test image. For text-based tasks (D1S and D2S), embeddings of image descriptions are obtained with fine-tuned MiniLM (HuggingFace, 2021; Wang et al., 2020a). Next, for each matrix side, the embedding that is most distant from the test embedding (based on Euclidean distance) is identified. The test input is then assigned to the side with the lower maximum distance (i.e., higher similarity). Appendix C.8 explains the reasons of SC strong performance.
@@
To better characterize model behavior, we categorized Bongard-RWR+ matrix concepts into $9$ semantic groups reflecting the key factor behind each concept: Size, Position, Count, Branching, Similarity, Contour, Shape, Rotation, and Angle. This grouping follows recent efforts advocating for semantically grounded evaluations to isolate model strengths and weaknesses (Mitchell, 2021; Odouard and Mitchell, 2022). The results are shown in Fig. 6. For $K=2$, all models exhibit broadly similar profiles across concept groups. However, for more challenging setups ($K>2$), differences emerge. For instance, InternVL2.5 retains high performance on Shape, Size, and Branching for $K=16$, with accuracy near $75\%$. These categories involve high-level visual attributes that VLMs appear to capture reliably. In contrast, InternVL2.5 accuracy drops below $50\%$ for Contour, Rotation, and Angle. These concepts often rely on subtle visual cues (Contour) or precise spatial relationships (Rotation and Angle), which current models struggle to capture reliably. We expand on this in Appendix C.10. Overall, performance patterns remain consistent across values of $K$, supporting the utility of such analysis as a diagnostic tool for characterizing model capabilities.
@@
Given the difficulty of the introduced dataset, we hypothesized that decomposing the problem via an intermediate captioning step may improve model performance. To test this, we generated image descriptions using each of the $4$ VLMs and used these captions as input to the same set of models for solving the side prediction task. Since the prediction task is purely text-based, we also evaluated the DeepSeek-R1 70B (DS-R1) reasoning model (Guo et al., 2025), which does not support visual input ("N/A" for I1S and I2S in Table 2) but offers strong language understanding. As shown in Fig. 7, final accuracy in the D1S task varies depending on both the captioning (rows) and prediction (columns) models. Descriptions produced by InternVL2.5 yield the highest overall accuracy, with Qwen2-VL also producing relatively strong captions. Interestingly, captions from MiniCPM-o 2.6 lead to better results than those from LLaVA-Next, despite MiniCPM-o 2.6 being much smaller, highlighting its promising image-to-text capabilities. Among prediction models (columns), Qwen2-VL achieves the best average performance, followed by InternVL2.5, indicating relatively strong text-based reasoning abilities of both models. DeepSeek-R1, while known for strong performance on standard benchmarks, ranks third, revealing potential limitations in reasoning over image descriptions. Using captions generated by the best-performing model, InternVL2.5, we report D1S and D2S task performance in Table 2. Compared to I1S and I2S, most models perform above chance in these settings, confirming that the intermediate captioning step helps models better ground their predictions.
@@
Figure 8: Concept Selection sensitivity analysis. (a) Accuracy of models of varying sizes for $K=16$. (b) The impact of using grayscale images for $K\in\{2,4,8,16\}$. Differences w.r.t. color images are shown using lighter colors and are annotated in parentheses above the corresponding bars.
@@
Prior works have shown that VLM performance tends to scale with model size (Hoffmann et al., 2022; Kaplan et al., 2020). To examine whether this observation holds for AVR, we evaluated smaller model variants in the CS task on Bongard-RWR+. Specifically, we tested 8B, 26B, and 34B variants of InternVL2.5, 7B, 32B, and 72B versions of LLaVA-Next and 7B variant of Qwen2-VL. We excluded additional MiniCPM-o 2.6 variants, as the 8B model already showed weak performance and no larger variants are currently available. As shown in Fig. 8(a), accuracy generally increases with model size. Among the smallest configurations (7B/8B), InternVL2.5 performs best, followed by Qwen2-VL, mirroring the relation observed for their largest versions (70+B). Notably, both InternVL2.5 8B and Qwen2-VL 7B outperformed MiniCPM-o 2.6 8B, underscoring the generally stronger AVR capacity of these two model families.
@@
To assess the role of color in abstract reasoning, we evaluated all $4$ VLMs on Bongard-RWR+/GS, using the CS task. As shown in Fig. 8(b), model performance remains comparable, or even improves, on grayscale images. For instance, LLaVA-Next consistently achieves higher accuracy across all $K$ (e.g. +5+5 p.p. for $K=2$ concepts). This outcome aligns with the dataset’s design – the underlying concepts are derived from structural properties of black-and-white synthetic BPs, where color is not semantically relevant. In this context, color may act as a visual distractor that adds additional complexity, especially for lower-performing models.
Figure 9: Functional equivalence of generated images and real-world images. Performance trends in the CS task ($K\in\{2,4,8,16\}$) on Bongard-RWR and Bongard-RWR+/LP for $P=2,3$ panels.
@@
Since Bongard-RWR+ relies on generated images, a key question is whether such images are as effective as real ones for evaluating visual reasoning. To investigate this, we compared model performance on the introduced dataset with that on Bongard-RWR, which consists exclusively of real-world images. However, given that Bongard-RWR includes only $60$ instances, we constructed matrices based on smaller subsets of images per side to expand the dataset. For each BP in Bongard-RWR, we considered all pairs of $P$-element subsets from both matrix sides (analogously as described in Section 3.1), yielding $\binom{6}{P}^{2}$ new matrices per BP, i.e., $36$, $225$, $400$, $225$, $36$ matrices for $P=1,\ldots,5$, resp. We applied this method to $54$ Bongard-RWR BPs that were used to generate Bongard-RWR+, and then selected $P=2,3$ to ensure sufficient dataset size, and capped the number of sampled matrices per source BP at $100$, resulting in 5 4005\,400 matrices per $P$. We evaluated all $4$ VLMs in the CS task on these Bongard-RWR variants and compared the results to those on Bongard-RWR+/LP, which uses analogous matrix construction approach but with generated images. As shown in Fig. 9, model performance follows similar trends on both datasets – accuracy decreases consistently as $K$ increases, regardless of whether the images are real or generated. This supports the validity of our generation-based approach for capturing the challenges of AVR.
@@
As $P$ increases, models are presented with more demonstrations of the underlying concept. While this could enhance concept identification, it also increases the amount of visual information to process. Interestingly, Fig. 9 shows that InternVL2.5 and Qwen2-VL benefit from larger $P$ in the CS task, e.g., InternVL2.5’s accuracy improves by $2,3,6,8$ p.p. for $K=2,\ldots,16$, resp. when comparing the results on Bongard-RWR+/L2 with Bongard-RWR+/L3. In contrast, the performance of LLaVA-Next and MiniCPM-o 2.6 shows no consistent improvement, suggesting these models may struggle to leverage additional examples.
To explore this further, we evaluated models on Bongard-RWR+/LP with $P=2,\ldots,6$ using I1S and D1S tasks (Table 3). In I1S, InternVL2.5 and Qwen2-VL show a mild upward trend with increasing $P$, with Qwen2-VL achieving the accuracy of $57\%$ at $P=6$, and InternVL2.5 close behind at $55\%$. In D1S, accuracy generally improves with $P$, showing that the models can effectively utilize additional concept demonstrations in a text-based setting. Notably, DeepSeek-R1 outperforms others for every $P>2$ (and is on-par with InternVL2.5 for $P=2$), demonstrating its strength in this setting. Additionally, D1S scores are consistently higher than I1S scores, showing that the models benefit from the explicit image captioning step. Nevertheless, performance remains modest on both tasks, in I1S being clearly inferior to the SC baseline, pointing to a persistent AVR gap of the current models.
Finally, models’ accuracy on Bongard-RWR+/LP exceeds that on Bongard-RWR+ (e.g., InternVL2.5 scores $50\%$ and $57\%$ in I1S and D1S, resp. on Bongard-RWR+ (cf. Table 2), but reaches $55\%$ and $67\%$ on Bongard-RWR+/L6). This gap stems from the differences in dataset construction, which we analyze in detail in Appendix C.
Figure 10: Accuracy vs. image diversity (CS, K=16K\!=\!16).
Visual diversity facilitates concept recognition.
Concept diversity can be expressed through both a larger number of examples per side ($P$) and greater variation among the images themselves. To measure diversity in a complementary way, we computed the maximum cosine similarity between ViT-L/14 image embeddings within each side of a BP, then took the negative of the larger value across both sides as the BP’s diversity score. BPs were partitioned into three equally sized bins (Low, Medium, High diversity), and we analyzed model performance in the most challenging CS variant ($K=16$).
As shown in Fig. 10, performance is lowest on Low-diversity matrices (e.g., InternVL2.5 78B scores $0.50$), where images are similar and depict the concept in similar ways, making it harder to identify. In contrast, performance is highest on High-diversity matrices (e.g., InternVL2.5 78B scores $0.62$, +12+12 p.p. vs. Low-diversity), where the concept is expressed through more varied visual content, making it more apparent. This indicates that diversity helps disambiguate abstract concepts.
Noisy images hinder concept recognition.
To quantify the impact of manual filtering on model performance, we conducted a targeted experiment introducing controlled noise into the dataset. Specifically, we constructed Bongard-RWR+/L2-Impure, a variant of Bongard-RWR+/L2 in which each BP side contains one image that passed human filtering and one image that was rejected. This yields BP instances where a subset of images violate the intended concept, producing ambiguity that would arise if human verification was removed.
We evaluated all four models on this impure dataset in the CS task (see details in Appendix C.5). Across $K\in\{2,4,8,16\}$ levels, we observe mean absolute accuracy drops of 3.85-3.85, 8.60-8.60, 5.73-5.73, and 4.60-4.60 for MiniCPM-o 2.6 8B, InternVL2.5 78B, Qwen2-VL 72B, and LLaVA-Next 110B, resp. This consistent reduction in accuracy supports our hypothesis: including unfiltered images degrades dataset quality and harms model performance, demonstrating that the human review step meaningfully improves data quality.
@@
We tested model ability to generate free-form concept descriptions in the CG task. Models were prompted to describe the concept behind each matrix, using either raw image inputs or InternVL2.5’s captions. Responses were evaluated using standard NLP metrics: BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), $\text{ROUGE}_{L}$ (Lin, 2004), CIDEr (Lin, 2004), and BERTScore (Zhang* et al., 2020). The results, reported in Appendix C, show that the models consistently achieved low scores on all metrics, with no clear best-performing model. Manual inspection of selected outputs further confirmed that current models generally struggle to articulate the BP concepts, highlighting this task as a particularly challenging direction for future research.
@@ 5 Conclusions
We introduced Bongard-RWR+, a large-scale dataset for evaluating AVR capabilities using generated real-world-like images of abstract concepts from classical BPs. The dataset supports diverse BP task formulations, including multiclass concept selection, binary side classification, and free-form text generation. Our results show that while current VLMs demonstrate certain ability to identify coarse-grained concepts, they consistently struggle with fine-grained concept recognition. Although explicit image captioning and increased visual demonstrations can improve performance, even the strongest models fall short in more demanding setups, notably unconstrained textual answer generation.
@@
Our dataset generation pipeline still requires human oversight to ensure adherence to intended concepts; advancing generative modeling is needed to improve scalability. Our evaluation includes selected VLMs; broader community involvement is essential to benchmark additional methods on Bongard-RWR+. We extend a discussion on limitations and their mitigation in Appendix D.
Ethics statement.
We note that images generated by the T2I model can reflect social, cultural, or geographical biases. Although our data generation algorithm incorporates diversity safeguards, some biases may still be present in the dataset. To address this, we conducted a limited-scale bias audit, which we further discuss in Appendices D and E.1.
In addition, since our pipeline (Fig. 3) involves manual review, we recognize the potential for annotation-related bias. To mitigate this, each generated image was independently reviewed by two expert annotators, who verified whether the image faithfully expressed the intended concept. Inter-annotator agreement reached Cohen’s $\kappa=0.64$ (substantial agreement), with annotators agreeing on the same label in $77.3\%$ of cases. Disagreements were resolved through discussion, and images lacking consensus were discarded. Overall, $30.2\%$ of generated images were excluded through this filtering process. A more detailed description of the review process and an annotator agreement report, are provided in Appendix E.2.
All images in our dataset are generated with Flux.1-dev; none of them was taken from public image databases or libraries. The dataset was released under the CC BY 4.0 License, as this licensing is compatible with the terms of the T2I model used for image generation [^1] https://huggingface.co/black-forest-labs/FLUX.1-dev/blob/main/LICENSE.md. We imposed an additional usage restriction by adding the following note: “While Bongard-RWR+ is released under the CC BY 4.0 License, it is explicitly not intended for use in high-stakes human assessment contexts, including but not limited to standardized testing, educational admissions, employment screening, or cognitive evaluation”.
Reproducibility.
All datasets developed in this work are publicly available via the HuggingFace Datasets platform [^2] https://huggingface.co/collections/bongard-rwr-plus/bongard-rwr-plus-68d2e8b4d1337884f132b2e7. Code for dataset generation and experimental evaluation is available on GitHub [^3] https://github.com/bongard-rwr-plus/bongard-rwr-plus under the MIT License.
Acknowledgments
This research was carried out with the support of the Laboratory of Bioinformatics and Computational Genomics and the High Performance Computing Center of the Faculty of Mathematics and Information Science Warsaw University of Technology.
Appendix A Broader Impact
Prior work in cognitive psychology shows that cultural background can indeed shape reasoning strategies and outcomes (Helms-Lorenz et al., 2003), sometimes leading to qualitatively different solution approaches (Nisbett et al., 2008). In particular, studies on RPMs discussed the variation between ethnic groups within the United States (Raven, 2005). While little work has examined BPs in this context, related studies in concept learning and categorization suggest that cross-cultural differences can in fact influence how abstract visual concepts are formed (Rosch, 1973). Building on this line of work, we believe that our Bongard-RWR+ dataset, which uses real-world images rather than abstract patterns, can serve as a complementary resource for studying how multi-image reasoning generalizes across different cultures.
Appendix B Dataset Variants